OpenStudy (anonymous):

Solve for x: x + 2 1/2 = 7. Also please EXPLAIN the work. I keep getting trouble on these problems.

OpenStudy (anonymous):

is that 2 1/2 like 2.5 and not \(\large 2^{\frac{1}{2}} \)??

OpenStudy (anonymous):

It's not the 2.5 part, I mean the mixed number part.

OpenStudy (anonymous):

okay... mixed number.... i shuda said that....:)

OpenStudy (anonymous):

subtract 2 1/2 from both sides to get: x = 7 - 2 1/2 = 4 1/2

OpenStudy (anonymous):

that mixed number is "four and one half."

OpenStudy (anonymous):

But how did you get the answer? Please explain

OpenStudy (anonymous):

7-2 = 5 so 7-2 1/2 = 4 1/2

OpenStudy (anonymous):

oh you want to see the subtraction...?

OpenStudy (anonymous):

oh no, I'm good I understand now

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):